An employment agency specializing on temporary construction helps pays heavy equipment operators $132 per day and general laborers $97 per day. Of thirty people were hired and the payroll was $3680,how many heavy equipment operators were employed? How many laborers?
Answer by stanbon(75887)

An employment agency specializing on temporary construction help pays heavy equipment operators $132 per day and general laborers $97 per day. Of thirty people were hired and the payroll was $3680,how many heavy equipment operators were employed? How many laborers?
-------------------
Equations:
h + g = 30 helpers
132h + 97g = 3680
--------------------------
g = 30-h
132h + 97(30-h) = 3680
-------------------------
132h + 2910 - 97h = 3680
35h = 770
h = 22 (# of heavy equip. operators)
g = 30-h = 8 (# of gen laborers)
---------------
Cheers,
Stan H.
----------